The expression for the number of molecules in a Maxwellian gas having speeds in the range vto v+ dv is


dNv= 4πN(m÷2πkbT)^3/2 v^2exp[-(mv^2÷2kbT)]dv using this relation, obtain an expression for most probable speed

1
Expert's answer
2022-02-24T09:05:13-0500
"dN=4πN\left(\frac{m}{2πkT}\right)^{3/2} v^2\exp[-(mv^2/2kT)]dv"

"( v^2\exp[-(mv^2/2kT)])'=2v\exp[-(mv^2/2kT)]\\
-mv^3/kT\exp[-(mv^2/2kT)]=0"

"2=mv^2\/kT"

"v_{mp}=\sqrt{2kT/m}"